OpenStudy (anonymous):

An architect drew the blueprint for a new office building. He used a scale in which 1 inch represents 7.5 feet. The floor of an office in the building will have actual dimensions of 18 feet by 16 feet. What will be the dimensions on the blueprint?

OpenStudy (anonymous):

A 25.5 inches by 23.5 inches B 135 inches by 120 inches C 2.4 inches by 2.13 inches

jimthompson5910 (jim_thompson5910):

1/7.5 = x/18 1*18 = 7.5x 18 = 7.5x 7.5x = 18 x = 18/7.5 x = ???

OpenStudy (anonymous):

2.4

jimthompson5910 (jim_thompson5910):

good